OpenStudy (anonymous):

Simple Harmonic Motion Question When finding the displacement-time function given the velocity-displacement function using integration, why do we assume v is positive? EDIT: Question is in comments.

OpenStudy (anonymous):

v^2=n^2(A^2-x^2) v=n(A^2-x^2)^(1/2) Why isn't v =+-n(A^2-x^2)^(1/2)

OpenStudy (anonymous):

\[v^{2}=n ^{2}(A ^{2}-x ^{2})\] \[v=n \sqrt{a^{2}-x ^{2}}\]

OpenStudy (anonymous):

If you need even the direction of velocity, then choose the sign depending on whether x>0 or x<0.

OpenStudy (anonymous):

But if you're just trying to find the displacement-time function, wouldn't your final answer be +- [the function]? But that wouldn't really make sense

OpenStudy (anonymous):

Yeah, if we consider the +ve sign and integrate, we get a trigonometric function which automatically takes care of sign.

OpenStudy (anonymous):

But depending on the sign, after we integrate, it will change to either arcsin or arccos...How do we figure out which one it is

OpenStudy (anonymous):

There should be some information of position and time, which can help us finding the appropriate function.

OpenStudy (anonymous):

A particle is moving with simple harmonic motion of period pi seconds and maximum velocity 8m/s. If the particle started from rest at x=a, find a.

OpenStudy (anonymous):

I got up to the step v^2=64-4x^2 but what do I do when I squareroot both sides? Do I put a +- or assume positive (like my textbook does). But why do you assume positive?

OpenStudy (anonymous):

Why don't you post it in physics? I will reply there..

OpenStudy (anonymous):

In Australia, motion is in mathematics for some reason haha :)

OpenStudy (anonymous):

That is good. :) You can solve this way, Use, \[T = 2\pi /\omega \] where T-time period, w- angular velocity. Now, use the expression \[v = A \omega \cos wt \] For maximum velocity. Got it?

OpenStudy (anonymous):

hmm but the book says to integrate to get an inverse trig function, and then rearrange the equation so its like x=[trig function]. Then you substitute t=0 in to find what a is

OpenStudy (anonymous):

Oh, then consider the positive sign. the sin or cos will take care of the sign. anyway, the w above was angular frequency. sorry.

OpenStudy (anonymous):

okay! I think I kind of get it... Is it like you can use positive or negative, but whichever you use, the trig function will take care of it for you?

OpenStudy (anonymous):

in the question above though, you have to take v as negative though